how do i solve 3(x+1/3)+4(x+1/2)=0
Answer by sehar(13)

the quetion is quite simple
3(x+1/3)+4(x+1/2)=0
x+1 + 2(x+1)=0
x+1 + 2x+2=0
3x+3=0
x=-3/3
x=-1